OpenStudy (anonymous):

Find the length of the diameter of a sphere with a surface area of 804.25. help please in solving it

OpenStudy (anonymous):

do you know the formula for the surface area of a sphere?

OpenStudy (anonymous):

4 pi r squared

OpenStudy (anonymous):

good... A = (4pi)r^2 dividing (4pi) from both sides we have: A/(4pi) = r^2 still with me?

OpenStudy (anonymous):

yes

OpenStudy (nali):

do u still need help

OpenStudy (anonymous):

yes please nali

OpenStudy (anonymous):

i cant seem to get the answer

OpenStudy (anonymous):

wait.. we're not done..

OpenStudy (anonymous):

A/(4pi) = r^2 take the square root of both sides we get: sqrt(A/(4pi)) = r that expression is the radius... do you know how to get the diameter?

OpenStudy (anonymous):

the diameter is twice the radius right?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

what did you get?